# On the Convergence of Multi-objective Optimization under Generalized Smoothness

### v2

@@ Abstract
Multi-objective optimization (MOO) is receiving more attention in various fields such as multi-task learning. Recent works provide some effective algorithms with theoretical analysis but they are limited by the standard $L$-smooth or bounded-gradient assumptions, which are typically unsatisfactory for neural networks, such as recurrent neural networks (RNNs) and transformers. In this paper, we study a more general and realistic class of $\ell$-smooth loss functions, where $\ell$ is a general non-decreasing function of gradient norm. We develop two novel single-loop algorithms for $\ell$-smooth MOO problems, Generalized Smooth Multi-objective Gradient descent (GSMGrad) and its stochastic variant, Stochastic Generalized Smooth Multi-objective Gradient descent (SGSMGrad), which approximate the conflict-avoidant (CA) direction that maximizes the minimum improvement among objectives. We provide a comprehensive convergence analysis of both algorithms and show that they converge to an $\epsilon$-accurate Pareto stationary point with a guaranteed $\epsilon$-level average CA distance (i.e., the gap between the updating direction and the CA direction) over all iterations, where totally $\mathcal{O}(\epsilon^{-2})$ and $\mathcal{O}(\epsilon^{-4})$ samples are needed for deterministic and stochastic settings, respectively. Our algorithms can also guarantee a tighter $\epsilon$-level CA distance in each iteration using more samples. Moreover, we propose an efficient variant of GSMGrad named GSMGrad-FA using only constant-level time and space, while achieving the same performance guarantee as GSMGrad. Our experiments validate our theory and demonstrate the effectiveness of the proposed methods.
@@ 1.1 Our Contributions
Weakest assumptions in MOO. In this paper, we investigate the $\ell$-smooth assumption, where $\ell$ is a general non-decreasing function of gradient norm, and includes both the standard $L$-smooth and $(L_{0},L_{1})$-smooth assumptions as special cases. This assumption finds many applications, such as transformers [35, 11], distributionally robust optimization [19] and higher-order polynomial functions [9]. In addition, we do not make any bounded-gradient assumption, which is required in previous analysis to ensure the bounded multi-gradient approximation. To the best of our knowledge, this is the first work to investigate generalized smoothness in MOO problems.
@@ 1.2 Related Works
Table 1: Comparison for existing stochastic methods for MOO problems to obtain an $\epsilon$–accurate Pareto stationary point. Explanation on the upper footmarks: $1:$ (LS) indicates that the objectives are standard $L$-smooth while (GS) the objectives are generalized $\ell$-smooth defined in Definition 1. $2:$ (BF) shows that the bounded function value assumption is required and (BG) indicates an additional bounded assumption is required. $3:$ the method in [25] only works for convex objectives and it is unable to obtain an $\epsilon$–accurate Pareto stationary point for non-convex objectives studied in this paper.
@@ 2.1 Generalized smoothness
The standard $L$-smoothness condition is widely investigated in existing optimization studies [15, 16], which assumes a function $f:\mathcal{X}\to\mathbb{R}$ to be $L$-smooth if there exists a bounded constant $L$ such that for any $x,y\in\mathcal{X}$, $\|\nabla f(x)-\nabla f(y)\|\leq L\|x-y\|.$ Nevertheless, recent studies show that in the training of neural networks such as transformers [35, 11], distributionally robust optimization [19] and high-order polynomials functions [9], the standard $L$-smoothness assumption does not hold. Instead, a generalized $(L_{0},L_{1})$-smoothness assumption was observed and studied in the training of transformers in [35], which assumes that for any $x\in\mathcal{X}$, $\|\nabla^{2}f(x)\|\leq L_{0}+L_{1}\|\nabla f(x)\|.$ This assumption implies the Lipschitz constant is potentially unbounded and reduces to the $L$-smoothness if $L_{1}=0$. Later, a more generalized assumption was proposed and studied in [21]:
@@ 5.1 Deterministic setting
Let Assumptions 1 and 2 hold. Set $N\sim O(\epsilon^{-2})$, $\beta^{\prime}\leq\frac{1}{M^{2}},\rho\leq\min\left(\mathcal{O}(\epsilon^{2}),\mathcal{O}\left(\frac{1}{\alpha T}\right)\right),\beta\leq\mathcal{O}(\epsilon^{2}),\alpha\leq\min\left(\mathcal{O}(\beta),\mathcal{O}(\epsilon^{2}),\mathcal{O}\left(\frac{1}{\beta T}\right)\right)$ as constants, and $T\geq\max\left(\Theta\left(\frac{1}{\alpha\epsilon^{2}}\right),\Theta\left(\frac{1}{\beta\epsilon^{2}}\right)\right)$. We then have $\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}\|^{2}\leq\epsilon^{2}.$

### v3

@@ Abstract
Multi-objective optimization (MOO) is receiving more attention in various fields such as multi-task learning. Recent works provide some effective algorithms with theoretical analysis but they are limited by the standard $L$-smooth or bounded-gradient assumptions, which are typically unsatisfactory for neural networks, such as Long short-term memory (LSTM) models and transformers. In this paper, we study a more general and realistic class of $\ell$-smooth loss functions, where $\ell$ is a general non-decreasing function of gradient norm. We develop two novel single-loop algorithms for $\ell$-smooth MOO problems, Generalized Smooth Multi-objective Gradient descent (GSMGrad) and its stochastic variant, Stochastic Generalized Smooth Multi-objective Gradient descent (SGSMGrad), which approximate the conflict-avoidant (CA) direction that maximizes the minimum improvement among objectives. We provide a comprehensive convergence analysis of both algorithms and show that they converge to an $\epsilon$-accurate Pareto stationary point with a guaranteed $\epsilon$-level average CA distance (i.e., the gap between the updating direction and the CA direction) over all iterations, where totally $\mathcal{O}(\epsilon^{-2})$ and $\mathcal{O}(\epsilon^{-4})$ samples are needed for deterministic and stochastic settings, respectively. Our algorithms can also guarantee a tighter $\epsilon$-level CA distance in each iteration using more samples. Moreover, we propose an efficient variant of GSMGrad named GSMGrad-FA using only constant-level time and space, while achieving the same performance guarantee as GSMGrad. Our experiments validate our theory and demonstrate the effectiveness of the proposed methods.
@@ 1.1 Our Contributions
Weakest assumptions in MOO. In this paper, we investigate the $\ell$-smooth assumption, where $\ell$ is a general non-decreasing function of gradient norm, and includes both the standard $L$-smooth and $(L_{0},L_{1})$-smooth assumptions as special cases. This assumption finds many applications, such as LSTM models [35], transformers [11], distributionally robust optimization [19] and higher-order polynomial functions [9]. In addition, we do not make any bounded-gradient assumption, which is required in previous analysis to ensure the bounded multi-gradient approximation. To the best of our knowledge, this is the first work to investigate generalized smoothness in MOO problems.
@@ 1.2 Related Works
Table 1: Comparison for existing stochastic methods for MOO problems to obtain an $\epsilon$–accurate Pareto stationary point. Explanation on the upper footmarks: $1:$ (LS) indicates that the objectives are standard $L$-smooth while (GS) the objectives are generalized $\ell$-smooth as defined in Definition 1; $2:$ (BF) shows that the bounded function value assumption is required and (BG) shows that the bounded gradient assumption is required; $3:$ the analysis in [25] focuses on convex objective functions.
@@ 2.1 Generalized smoothness
The standard $L$-smoothness condition is widely investigated in existing optimization studies [15, 16], which assumes a function $f:\mathcal{X}\to\mathbb{R}$ to be $L$-smooth if there exists a bounded constant $L$ such that for any $x,y\in\mathcal{X}$, $\|\nabla f(x)-\nabla f(y)\|\leq L\|x-y\|.$ Nevertheless, recent studies show that in the training of neural networks such as LSTM models [35], transformers [11], distributionally robust optimization [19] and high-order polynomials functions [9], the standard $L$-smoothness assumption does not hold. Instead, a generalized $(L_{0},L_{1})$-smoothness assumption was observed and studied in the training of LSTM models in [35], which assumes that for any $x\in\mathcal{X}$, $\|\nabla^{2}f(x)\|\leq L_{0}+L_{1}\|\nabla f(x)\|.$ This assumption implies the Lipschitz constant is potentially unbounded and reduces to the $L$-smoothness if $L_{1}=0$. Later, a more generalized assumption was proposed and studied in [21]:
@@ 5.1 Deterministic setting
Let Assumptions 1 and 2 hold. Set $N\sim O(\epsilon^{-2})$, $\beta^{\prime}\leq\frac{1}{M^{2}},\rho\leq\min\left(\mathcal{O}(\epsilon^{2}),\mathcal{O}\left(\frac{1}{\alpha T}\right)\right),\beta\leq\mathcal{O}(\epsilon^{2}),\alpha\leq\min\left(\mathcal{O}(\beta),\mathcal{O}(\epsilon^{2}),\mathcal{O}\left(\frac{1}{\beta T}\right)\right)$ as constants, $T\geq\max\left(\Theta\left(\frac{1}{\alpha\epsilon^{2}}\right),\Theta\left(\frac{1}{\beta\epsilon^{2}}\right)\right)$. We have $\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}\|^{2}\leq\epsilon^{2}.$